### v4

# From Informal to Formal – Incorporating and Evaluating LLMs on Natural Language Requirements to Verifiable Formal Proofs

Abstract
The research in AI-based formal mathematical reasoning has shown an unstoppable growth trend. These studies have excelled in mathematical competitions like IMO and have made significant progress. However, these studies intertwined multiple skills simultaneously—problem-solving, reasoning, and writing formal specifications—making it hard to precisely identify the LLMs’ strengths and weaknesses in each task. This paper focuses on formal verification, an immediate application scenario of formal reasoning, and breaks it down into sub-tasks. We constructed 18k high-quality instruction-response pairs across five mainstream formal specification languages (Coq, Lean4, Dafny, ACSL, and TLA+) in six tasks by distilling gpt-4o and evaluated against ten open-sourced LLMs, including recent popular DeepSeek-R1. We found that LLMs are good at writing proof segments when given either the code, or the detailed description of proof steps. Also, the fine-tuning brought about a nearly threefold improvement at most. Interestingly, we observed that fine-tuning with formal data also enhances mathematics, reasoning, and coding capabilities. Fine-tuned models are released to facilitate subsequent studies at https://huggingface.co/fm-universe.
From Informal to Formal – Incorporating and Evaluating LLMs on
Natural Language Requirements to Verifiable Formal Proofs
Jialun Cao1, Yaojie Lu2, Meiziniu Li1, Haoyang Ma1, Haokun Li3, Mengda He3, Cheng Wen3, Le Sun2, Hongyu Zhang4, Shengchao Qin3,5, Shing-Chi Cheung1, Cong Tian5 1 The Hong Kong University of Science and Technology 2 Institute of Software, Chinese Academy of Sciences 3 Guangzhou Institute of Technology, Xidian University 4 Chongqing University 5 ICTT and ISN Laboratory, Xidian University jcaoap@cse.ust.hk, luyaojie@iscas.ac.cn
1 Introduction
The more we formalize, the more of our implicit knowledge becomes explicit. ❞ — Terence Tao Tao (2024)
Figure 1: The Illustration of Formal Proof Generation and Its Relation with Code Generation.
As AI-based formal mathematical reasoning reached an inflection point Yang et al. (2024b), significant attention and progress in this field have been observed. AlphaProof AlphaProof and Teams (2024) achieved silver medal level in the International Mathematical Olympiad (IMO), AlphaGeometry Trinh et al. (2024) specialized in proving Euclidean geometry theorems. As reported, the number of publications in this field nearly doubled in 2023, indicating an unstoppable growth trend Li et al. (2024). As Fields Medalist Terence Tao imagined, “In the future, instead of typing up our proofs, we would explain them to some GPT” Tao (2024).
However, most current benchmarks cannot precisely reflect the capability to convert informal proofs or requirements in natural language into formal proofs. Most of these benchmarks take mathematical problems AlphaProof and Teams (2024); Trinh et al. (2024); Welleck et al. (2021) or theorems to be solved Yang et al. (2023); Welleck et al. (2022); Yang and Deng (2019) as input, and informal or formal proofs (or parts of proofs) as output. However, these end-to-end benchmarks assess multiple capabilities (e.g., problem-solving, mathematical reasoning, formal specification writing) in an intertwined manner, making it difficult to isolate and observe LLMs’ true capabilities in writing formal proofs or models for verification.
Therefore, we break down the process from informal requirements to formal verifiable proof, as shown in Figure 1. Inspired by the code generation (shown in blue) which translates a description of implementation into executable code Chen et al. (2021); Austin et al. (2021a), the formal reasoning process (shown in green) can be seen as translating an informal requirement into a verifiable formal proof or checkable formal model  [^1] For ease of expression, we generally refer to verifiable formal proofs and checkable models as “proofs” for the sake of presentation simplification.. Particularly, we decompose this process and formulate six tasks (Figure 2). By doing so, the intertwined capabilities can be separated and individually assessed, providing a clearer understanding of LLMs’ strengths and weaknesses in each task.
Scope and Targets – We focus on formal verification Appel (2011); Klein et al. (2009); Leroy et al. (2016); Hawblitzel et al. (2014) because it is an immediate application scenario of formal mathematical reasoning and the correctness of the output can be verified mechanically. In this paper, we mainly explore four research questions (RQs):
RQ1. How well do LLMs perform in various formal verification tasks? After decomposing the formal verification task into subtasks, we explore LLMs’ initial performance in these tasks with zero-shot and few-shot, investigating the strengths and weaknesses that vary between LLMs and tasks.
RQ2. Do LLMs show variability in their capability across different formal specification languages? When mathematicians and proof engineers consider using LLMs to assist in formal verification, they often face uncertainty about which formal specification language is best supported by LLMs. This RQ is designed to provide hints on it.
RQ3. Can fine-tuning improve LLMs’ performance in formal verification? Although recent efforts have been made to fine-tune models Wang et al. (2024); Yang et al. (2023), these LLMs are typically fine-tuned with single formal languages instead of multi-lingual (e.g., combining Coq, Lean, etc.) Yang et al. (2024b). Therefore, we instruction fine-tuned Wei et al. (2021); Sanh et al. (2021) three base LLMs to see whether our constructed fine-tuning dataset fm-alpaca could improve their capability in formal verification tasks.
RQ4. Can fine-tuning with formal verification data benefit other related tasks (mathematics, reasoning, code)? As recent works have shown LLMs’ potential transferability of skills Tihanyi et al. (2023) we thus extend our study to see if models fine-tuned on formal data could show enhanced capabilities in mathematics, reasoning, and coding.
To facilitate the study, we constructed 18k high-quality instruction-response pairs across five formal specification languages (i.e., Coq, Lean4, Dafny, ACSL, and TLA+) in six formal-verification-related tasks by distilling gpt-4o inspired by prior work Wang et al. (2024); Ding et al. (2023); Wang et al. (2022), then split them into 14k instruction fine-tuning data (fm-alpaca) and 4k benchmarking data (fm-bench). In particular, we provide executable contexts for all these formal specifications and automated validation scripts to validate the correctness of the generated formal proofs inspired from the prior work’s artifact preparation Jimenez et al. (2023). Finally, we release the fine-tuned LLMs based on three base models at https://huggingface.co/fm-universe.
Interestingly, there has been recent discussion on the topic of domain transfer Yang et al. (2024b), particularly the transfer of knowledge from other domains such as coding and reasoning to formal domains in order to increase LLMs’ reliability Spiess et al. (2024), and the anticipated potential of AI in enhancing formal verification processes to support mathematical proofs Tao (2024); Tao et al. (2023). Our experimental results could potentially provide empirical support for these hypotheses or offer directions for further experimental inquiries.
The contribution of this paper includes:
Problem Formulation: We decompose the formal verification process into six essential tasks. By doing so, the intertwined capabilities can be separated and individually assessed, providing a clearer understanding of LLMs’ strengths and weaknesses in each task.
Dataset and Benchmark: We constructed 18k high-quality instruction-response pairs across five mainstream formal specification languages (i.e., Coq, Lean4, Dafny, ACSL, and TLA+) in six formal-verification-related tasks by distilling gpt-4o. They are split into a 14k+ fine-tuning dataset fm-alpaca and a 4k benchmark fm-bench.
Executable context and automated validation mechanism: We provide a Docker container equipped with necessary scripts to facilitate the evaluation of fm-bench, significantly lowering the entry barrier for this scenario and making subsequent contributions easier.
Insight and Vision: We fine-tuned several models on fm-alpaca and observed promising benefits to not only the formal verification tasks, but also mathematics, reasoning, and coding. Our experimental results provide empirical support for the potential of LLMs’ capability transfer and hope to shed some light on future research.
2 Task Formulation
Figure 2 illustrates the six sub-tasks. We elaborate on them in detail as follows.
Task 1. Requirement Analysis (abbrev. ReqAna). Requirement analysis Davis (1990); Anton (1996); Grady (2010); Jin (2017) is a critical and long-standing research area in software engineering. It facilitates collecting, identifying, categorizing and modeling the users’ needs and expectations using various techniques Taggart Jr and Tharp (1977); Deeptimahanti and Babar (2009); Javed and Lin (2021); Wang et al. (2021); Jin et al. (2024); Zhou et al. (2022). In this paper, the requirements are the descriptions in natural language (English) Jin et al. (2024) that details the requirements of the verification/modeling goal and an overall description of the proofs/models. The task is to analyze and break down the final goal into detailed steps described in natural language. The natural language used in this paper is English.
Task 2. Full Proof Generation (abbrev. ProofGen). This task formalizes a requirement in natural language into verifiable proofs or models written in formal specification languages, similar task formulation to existing works Fatwanto (2012); Zhou et al. (2022); Davril et al. (2013).
Task 3. Proof Segment Generation (abbrev. SegGen). Unlike ProofGen, which requires generating complete proofs/models, SegGen provides more detailed descriptions in natural language and requires LLMs to write less. Given a text description articulating how to implement the proofs/modeling, the task outputs a segment written in the formal specification that serves as a component in the complete proof/model. This task formulation is similar to prior work Wang et al. (2024); Wu et al. (2022); Jiang et al. (2022) and similar to the formulation of code generation Chaudhary (2023); Sun et al. (2024); Welleck et al. (2022, 2021).
Task 4. Proof Completion (abbrev. ProofComplete). Similar to code completion Raychev et al. (2014); Husein et al. (2024); Svyatkovskiy et al. (2019); Dakhel et al. (2023), ProofComplete suggests the suffix of the given prefix, similar to prior work Song et al. (2024). Note that in order to prevent LLMs from deviating from the original verification goal, we also provide the requirement in our evaluation, although it is not compulsory for this task formulation.
Task 5. Proof InFilling (abbrev. ProofInfill). Given a proof/model with a mask in the middle, the task requires LLMs to fill proper formal specifications so that the completed proofs/models can pass the verifier. This formulation is the same as code infilling Fried et al. (2022). Also, similar to ProofComplete, we provide the requirement in our evaluation during the infilling to prevent LLMs from deviating from the original verification goal.
Task 6. Proof Generation from Code (abbrev. Code2Proof). In addition to generating formal specifications from natural languages, formal specifications can also be generated from code if the verification goal is the property of a given program. In this paper, we focus mainly on specifications in form of code annotations Baudin et al. (2021); Hatcliff et al. (2012), expressing specifications (e.g., pre-/post-condition, loop invariants) that help one to verify that (part of) a program satisfies certain properties. The task takes the code with properties to be verified as input and outputs the code with generated annotated formal specifications. Similar task formulation can be found in recent works Wen et al. (2024); Ma et al. (2024).
Figure 2: Six tasks towards Informal to Formal Verification
3 Data Construction
3.1 Formal Specification Language Selection
In this study, we consider five formal specification languages that can be used for formal verification, including Coq Huet (1986), Dafny Leino (2010), Lean4 Moura and Ullrich (2021), ACSL (ANSI/ISO C Specification) Baudin et al. (2021) and TLA+ Yu et al. (1999); Lamport (2002). We selected them in order to cover various verification paradigms (i.e., theorem proving and model checking) and verification scenarios (e.g., mathematical reasoning and program verification).
First, for interactive theorem provers which are suitable for developing rigorous mathematical proofs, we consider Coq Huet (1986) and Lean4 Moura and Ullrich (2021) because Coq has been extensively used in academia and research for proving mathematical theorems and in formal verification of software for a long history, while Lean4 garnered considerable attention from the mathematical community Wang et al. (2024); Tao et al. (2023); Avigad et al. (2020) recently. Second, for programming languages with built-in specification, we consider Dafny Leino (2010) and ACSL Baudin et al. (2021); Cuoq et al. (2012) because they seamlessly integrate specifications (e.g., pre-/post-conditions, loop invariants) within the code, ensuring the correctness through embedded assertions and conditions. Lastly, for model checking Jhala and Majumdar (2009); Clarke (1997), we consider TLA+ Yu et al. (1999) since it is a representative math-based formal language for modeling algorithms and programs such as concurrent and distributed systems.
3.2 Data Preparation
Figure 3: The Illustration of Data Preparation.
The workflow of data preparation for fm-alpaca and fm-bench is illustrated in Figure 3. The workflow begins with the data collection, where formal proofs in the desired formal specification languages and related configurations and dependencies are gathered from open-source repositories in Github. Then, formal proofs are extracted from the collected repositories. Next, the proofs go through the data quality assurance check by execution, the proofs that cannot be verified successfully are filtered out. The remaining ones are split into segments (e.g., definition of functions or conditions).
Given the impracticality of manually writing descriptions for all the collected formal proofs, we leveraged distilled GPT4 gpt (2023) to generate high-quality informal proof descriptions via meticulous prompting. This alternative is well-established and frequently employed in prior literature Wang et al. (2022, 2024); Ding et al. (2023); Wang et al. (2023). Specifically, for each formal specification language, we designated the model as an expert in that particular language, equipping it with comprehensive domain knowledge about the language’s specifications, essential grammatical cues, and three-shot examples featuring proof segments in the formal language as inputs and natural language descriptions as outputs. This approach ensures that the collected descriptions are of high quality and well-organized. It’s important to note that we did not generate descriptions for proof segments shorter than two lines (such as package imports or constant definitions) because their meaning is self-explained, with the exception of ACSL, whose proof segments are typically 1-2 lines.
After the descriptions for both full and segment proofs were prepared, we then prepared the data pairs for each task as shown in Task-wise Data Pairing in Figure 3. Note that for Task 4, i.e., Proof Completion, to prepare the incomplete formal proof, we randomly choose a line number and delete the lines in the proof after the line. For Task 5, i.e., Proof Infill, we randomly choose two line numbers and mask the lines between them. In case the remaining lines of proof cannot provide sufficient information for the proof generation, we also provide the informal proof for these two tasks.
After pairing the instruction-response for different tasks, we manually designed five task instructions for each task and randomly assigned one for each paired data to increase instruction diversity and avoid overfitting to certain instructions Lu et al. (2022); Feng et al. (2023); Sanh et al. (2022).
Spec Num of Proofs Num of Segments
Total fm-alpaca fm-bench Total fm-alpaca fm-bench
Coq 2,126 1,683 443 14,939 11,638 3,301
Lean4 1,163 919 244 1,578 1,261 317
ACSL 544 426 118 765 598 167
Dafny 249 206 43 417 348 69
TLA+ 256 199 57 594 476 118
Total 4,338 3,433 905 18,293 14,321 3,972
Average 868 687 181 3,659 2,864 794
Table 1: Formal-Specification-Language-wise Statistics of Formal Verification Data
3.3 Data Statistics
[…]
Task Total fm-alpaca fm-bench
1 Requirement Analysis 627 496 131
2 Full Proof Generation 700 557 143
3 Segment Proof Generation 14843 11597 3246
4 Proof Complete 658 520 138
5 Proof Infill 1439 1146 293
6 Code2Proof 70 56 14
Total 18337 14372 3965
Average 3056 2395 661
Table 2: Task-wise Statistics of Formal Verification Data
Table 2 shows the task-wise statistics. There are 18k instructions across six tasks, with an average of 3k instructions for each task. After splitting the train-test set, fm-alpaca contains 14k, and fm-bench has nearly 4k instructions. It is clear that the number of instructions for the task Segment Proof Generation (SegGen) is far more than that for Requirement Analysis (ReqAna) and Full Proof Generation (ProofGen) because one full proof can be split into numerous pieces of proof segments, and one proof can contribute to only one instruction for ReqAna and ProofGen. Note that the number of ReqAna (627) and ProofGen (700) is unequal because we filtered out the instructions with more than 2048 tokens considering the context limits.
3.4 Validation Mechanism
For the tasks whose outputs are written in formal specification languages, we verify the full proofs against the corresponding verifiers, i.e., Coq, Dafny, Lean4 use their own proving environment; formal specification written in TLA+ can be checked by TLC Yu and Kuppe ; C programs with ACSL specifications can be checked by Frama-C Cuoq et al. (2012); Carvalho et al. (2014). Also, for the proof segments that cannot be verified independently, for each extracted segment, we prepared a proof template with a placeholder during data preparation. Whenever a generated segment is to be verified, we replace the placeholder in the template with the segment and verify the completed formal proof.
For ACSL specification verification, we report the results under two modes. First, Basic checking. ACSL uses the ‘frama-c’ tool for syntax checking (i.e., , whether the syntax is correct) and other basic ACSL checking. Second, Verification which involves semantics checking (i.e., program behaviors are consistent with the annotated ACSL specifications) and verification towards the property under verification. ‘ACSL-WP’ uses ‘frama-c’ with the WP plugin for formal verification. This mode requires not only the basic checking (i.e., , syntactically correct, semantically consistent with the program), but also that the properties under verification can be proved.
Note that the invalid outputs (e.g., empty response or responses without formal specifications) are considered incorrect directly and will not go through the verification.
For the task whose outputs are written in natural language (i.e., ReqAna), we calculate the Bleu score Papineni et al. (2002) between the descriptions in fm-bench with the predicted outputs.
ProofGen SegGen ProofComplete ProofInfill Cd2Prf
LLMs Size TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL-WP ACSL ACSL-WP Ave
W/o Fine-tuning
llama3.1-instruct 8B 0.00 4.26 1.69 0.00 14.29 0.00 1.43 1.19 6.06 8.33 0.00 0.00 0.00 3.57 5.56 28.57 0.00 4.85 7.69 13.33 21.43 0.00 57.14 0.00 7.47
llama3.1-instruct 70B 7.69 6.38 10.17 20.00 7.14 0.00 24.29 4.86 13.64 11.11 39.68 3.17 27.27 3.57 13.89 28.57 0.00 7.27 24.62 13.33 21.43 0.00 21.43 0.00 12.90
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 21.43 0.00 1.43 1.05 7.58 5.56 1.59 1.59 0.00 13.10 19.44 28.57 0.00 5.45 18.46 13.33 21.43 0.00 14.29 0.00 7.56
qwen2.5-instruct 72B 7.69 6.38 18.64 20.00 21.43 7.14 12.86 3.29 12.12 25.00 12.70 4.76 0.00 11.90 33.33 28.57 0.00 9.70 24.62 26.67 35.71 0.00 57.14 0.00 15.82
qwen2.5-coder-instruct 7B 0.00 2.13 6.78 0.00 0.00 0.00 2.86 2.14 12.12 11.11 3.17 0.00 0.00 4.76 11.11 14.29 0.00 3.64 23.08 20.00 0.00 0.00 0.00 0.00 4.88
qwen2.5-coder-instruct 32B 0.00 6.38 11.86 0.00 21.43 0.00 12.86 4.07 9.85 25.00 26.98 4.76 27.27 5.95 36.11 42.86 2.94 7.88 33.85 20.00 71.43 14.29 57.14 14.29 19.05
starcoder-instruct 15B 7.69 2.13 8.47 30.00 28.57 0.00 27.14 1.97 6.06 27.78 0.00 0.00 27.27 11.90 25.00 42.86 0.00 9.70 24.62 20.00 35.71 0.00 28.57 0.00 15.23
deepseek-coder-instruct 7B 0.00 2.13 8.47 0.00 0.00 0.00 4.29 1.73 3.79 5.56 0.00 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 0.00 0.00 0.00 0.00 5.42
deepseek-coder-instruct 33B 0.00 0.00 1.69 0.00 0.00 0.00 2.86 1.94 4.55 11.11 3.17 0.00 0.00 3.57 13.89 42.86 0.00 2.42 23.08 6.67 7.14 0.00 7.14 0.00 5.50
deepseek-r1 671B 30.77 8.51 30.51 30.00 35.71 7.14 22.86 10.70 21.21 22.22 49.21 3.17 45.45 15.48 33.33 28.57 5.88 9.70 24.62 6.67 42.86 21.43 64.29 14.29 24.36
Task-wise average 7.70 9.16 18.31 12.62 16.79
W/ Fine-tuning
llama3.1-fma 8B 0.00 6.38 8.47 20.00 0.00 0.00 38.57 29.58 21.97 25.00 88.89 17.46 36.36 8.33 11.11 28.57 0.00 0.00 7.69 0.00 21.43 0.00 21.43 0.00 16.30
llama3.1-ultrachat 8B 0.00 6.38 3.39 0.00 0.00 0.00 0.00 0.00 3.79 0.00 0.00 0.00 0.00 5.95 13.89 14.29 0.00 4.85 9.23 0.00 14.29 7.14 0.00 0.00 3.47
llama3.1-ultrachat-fma 8B 0.00 8.51 10.17 30.00 0.00 0.00 41.43 35.79 29.55 33.33 95.24 17.46 18.18 8.33 11.11 28.57 2.94 0.00 7.69 0.00 28.57 0.00 35.71 0.00 18.44
llama3.1-tulu 8B 0.00 2.13 1.69 0.00 0.00 0.00 0.00 0.71 6.06 0.00 0.00 0.00 0.00 5.95 8.33 28.57 0.00 1.82 10.77 0.00 7.14 0.00 14.29 0.00 3.64
llama3.1-tulu-fma 8B 0.00 4.26 11.86 30.00 7.14 0.00 42.86 36.64 27.27 36.11 98.41 17.46 18.18 10.71 11.11 42.86 2.94 2.42 6.15 0.00 35.71 0.00 50.00 0.00 20.50
qwen2.5-fma 7B 0.00 4.26 11.86 10.00 0.00 0.00 38.57 27.95 27.27 22.22 87.30 17.46 27.27 9.52 13.89 28.57 0.00 0.00 7.69 0.00 21.43 0.00 14.29 0.00 15.40
qwen2.5-coder-fma 7B 0.00 6.38 18.64 20.00 7.14 0.00 44.29 36.50 34.09 33.33 98.41 17.46 36.36 9.52 16.67 28.57 5.88 1.82 13.85 13.33 42.86 0.00 42.86 0.00 22.00
deepseek-coder-fma 7B 0.00 2.13 16.95 0.00 0.00 0.00 34.29 25.30 31.06 25.00 84.13 15.87 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 14.29 7.14 14.29 7.14 16.10
Task-wise average 6.50 39.82 $\uparrow$ 20.69 $\uparrow$ 7.22 15.48
* -fma: fine-tuned with fm-alpaca. * -ultrachat: fine-tuned with UltraChat. * -tulu: fine-tuned with Tulu3. * -ultrachat-fma: fine-tuned with both UltraChat and fm-alpaca. * -tulu-fma: fine-tuned with both Tulu3 and fm-alpaca.
* ACSL and ACSL-WP: ‘ACSL’ uses ‘frama-c’ tool for syntax checking and other basic ACSL checking; ‘ACSL-WP’ uses ‘frama-c’ with the WP plugin for specification verification.
The task-wise average of finetuned models is calculated excluding the models (i.e., llama3.1-ultrachat and llama3.1-tulu) that are finetuned without fm-alpaca to remove the influence of the average value of the control group models’ results.
Table 3: RQ1-3: Pass@1 Accuracy of LLMs’ Performance Across Formal Verification Task and Formal Specification Languages with (w/) and without (w/o) fine-tuning. The greener, the better.
4 Experiments
4.1 Experiment Setup
Studied LLMs. We selected ten LLMs as baselines without fine-tuning, including llama3.1-instruct-8B/70B (Meta, 2024), qwen2.5-instruct-7B/72B (Yang et al., 2024a), qwen2.5-coder-instruct-7B/-32B Hui et al. (2024), starcoder-instruct-15B Lozhkov et al. (2024), deepseek-coder-instruct-7B-v1.5, deepseek-coder-instruct-33B (Guo et al., 2024), and deepseek-R1 Guo et al. (2025). Note that we avoid evaluating the GPT-series LLMs by OpenAI because the descriptions in fm-bench were generated by GPT-4o, making the evaluation fairer.
Fine-tuning. Instruction fine-tuning Wei et al. (2021); Sanh et al. (2021); Ding et al. (2023); Ivison et al. (2023) aims to improve a model’s ability to effectively respond to human instructions and has shown strong experimental potential in model enhancement. We select llama3.1-8B-instruct Meta (2024), qwen2.5-7B-instruct Yang et al. (2024a), and deepseek-coder-7B-instruct-v1.5  Guo et al. (2024) as base models for fine-tuning. These three models are chosen due to their promising performance on tasks such as coding, mathematics, and reasoning, and fine-tuning models in their scale is relatively affordable compared with fine-tuning larger scale models. We fine-tuned the three aforementioned models over three epochs using a learning rate 2e-5, a warm-up ratio of 0.04, a batch size of 512, and a cosine learning rate scheduler.
ReqSplitter
LLMs Size TLA Coq Lean Dafny ACSL
w/o fine-tune
llama3.1-instruct 8B 0.33 0.26 0.35 0.36 0.28
llama3.1-instruct 70B 0.33 0.28 0.35 0.41 0.30
qwen2.5-instruct 7B 0.29 0.25 0.32 0.33 0.19
qwen2.5-instruct 72B 0.33 0.28 0.33 0.34 0.22
qwen2.5-coder-instruct 7B 0.29 0.26 0.35 0.42 0.25
qwen2.5-coder-instruct 32B 0.31 0.27 0.33 0.39 0.23
star-coder-instruct 15B 0.21 0.24 0.29 0.30 0.29
deepseek-coder-instruct 7B 0.34 0.30 0.39 0.49 0.36
deepseek-coder-instruct 33B 0.37 0.29 0.47 0.47 0.43
deepseek-r1 671B 0.32 0.27 0.33 0.27 0.28
Average 31.85
w/ fine-tune
llama3.1-fma 8B 0.33 0.18 0.49 0.49 0.42
llama3.1-ultrachat 8B 0.35 0.30 0.44 0.40 0.46
llama3.1-ultrachat-fma 8B 0.37 0.30 0.56 0.57 0.62
llama3.1-tulu 8B 0.41 0.30 0.43 0.42 0.39
llama3.1-tulu-fma 8B 0.57 0.33 0.57 0.59 0.71
qwen2.5-fma 7B 0.33 0.13 0.53 0.58 0.38
qwen2.5-coder-fma 7B 0.51 0.21 0.57 0.64 0.73
deepseek-coder-fma 7B 0.43 0.24 0.46 0.54 0.43
Average 44.36 (39% $\uparrow$)
* -fma: fine-tuned with fm-alpaca.
Table 4: Evaluation on Requirement Analysis
Baseline Fine-tuning Datasets: To distinguish whether the capability improvement is simply because more instruction tuning is applied, we also include two commonly used fine-tuning datasets for comparison. We select UltraChat Ding et al. (2023) and Tulu-V3 Lambert et al. (2024) as baseline fine-tuning datasets, and use llama3.1–8B-base Meta (2024) as the base model due to their popularity. In particular, UltraChat is a large-scale dataset of instructional conversations that contains 1.5 million high-quality multi-turn dialogues and covers a wide range of topics and instructions. Tulu-v3 Lambert et al. (2024) embraces new data that is either carefully manu ally curated for quality or generated from GPT models. It is an enhancement of its previous versions Ivison et al. (2023); Wang et al. (2023), focusing more on core skills of knowledge recall, reasoning, mathematics, coding, instruction following, general chat, and safety.
Benchmarks for Related Capabilities (RQ4). To comprehensively evaluate the model’s capabilities, we tested the fine-tuned models on a series of benchmarks: Math (Hendrycks et al., 2021) and GSM-8K (Cobbe et al., 2021) for mathematical reasoning, BBH (Suzgun et al., 2022) for general reasoning, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021b) for coding.
Inference Strategies: We adopt different settings for different RQs. In particular, We use (1) the greedy sampling strategy to generate one single greedy sample with a temperature of 0.0 and calculate Pass@1, and (2) nucleus sampling Holtzman et al. (2020), where five solution samples were randomly generated with a temperature of 0.2 for RQ1 and RQ2. We also consider different in-context learning strategies, including zero-shot and few-shot (we used 3-shot in the experiment). For RQ3, we use a zero-shot greedy search with a temperature of 0.0 and a few-shot nuclear search with a temperature of 0.2 for a fair comparison.
Experiment Environment. The fine-tuning experiment was conducted on 32 Nvidia A800-40G GPUs, while inference was on Nvidia A100-80G GPUs with vLLM (Kwon et al., 2023).
4.2 RQ1. Basic Performance across Formal Specification Tasks
To understand the current LLMs’ performance in six tasks, we evaluate 10 LLMs against fm-bench with model size ranges from 7B to 671B. The upper part of Table 3 and the upper part of Table 4 show LLMs’ basic performance without fine-tuning.
Task-wise: LLMs perform the best in ProofComplete (18.31%) and Code2Proof (16.79%). In contrast, LLMs fall short in generating both the entire formal proof (7.70%) and the proof segments (9.16%). We analyzed the failures and found that syntax errors account for a large proportion, with 12.43% failures caused by syntax errors (See Appendix B). The observation echoes the motivation of prior work Wang et al. (2024) and is reasonable due to the grammar difference between most formal specification languages and other programming languages like Python. Regarding requirement analysis, as shown in the upper part of Table 4, the Bleu scores between the ground-truth description and LLM-generated ones range from 0.19 to 0.49.
LLM-wise: Without fine-tuning, DeepSeek-R1 achieved the best average (24.36%), followed by qwen2.5-coder-instruct-32B (19.05%).
Model Size: Larger LLMs generally perform better than smaller LLMs. For example, llama3.1-8B only achieved 1.43% in generating TLA+ segments, while llama3.1-70B boosts to 24.29% in the same task. However, there are several exceptions worth noticing, especially for ProofInfill and Code2Proof. For example, llama-3.1-8B achieved 57.14% in Code2Proof (ACSL), yet the performance drops to 21.43% using the llama-3.1-70B model. The decrease in performance is inherently due to the fine-tuning strategy of these instruction models: they are trained to excel in generating rather than filling in the blanks Fried et al. (2022). Also, we conducted a more detailed examination of generated segments and observed that larger LLMs tend to fill in the proof segments that not only look more plausibly correct and well-organized but also include extra content. The additional content, yet, is either redundant, as it repeats information that appears in the subsequent proof, or is incomplete. Promisingly, recent model developers have noticed such conundrums and refined their fine-tuning strategy for fill-in-the-middle tasks Guo et al. (2024).
TLA Coq Lean Dafny ACSL ACSL-WP
Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot
LLMs Size P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1
llama3.1-instruct 8B 1.43 4.29 11.43 1.19 1.90 10.12 6.06 8.33 9.85 8.33 11.11 19.44 0.00 3.17 38.10 0.00 0.00 12.70
llama3.1-instruct 70B 24.29 27.14 38.57 4.86 6.01 3.46 13.64 17.42 15.91 11.11 13.89 16.67 39.68 52.38 76.19 3.17 4.76 14.29
qwen2.5-instruct 7B 1.43 2.86 5.71 1.05 1.39 12.80 7.58 9.09 9.85 5.56 5.56 13.89 1.59 3.17 66.67 1.59 3.17 15.87
qwen2.5-instruct 72B 12.86 17.14 24.29 3.29 3.80 21.05 12.12 14.39 19.70 25.00 25.00 25.00 12.70 15.87 90.48 4.76 4.76 15.87
qwen2.5-coder-instruct 7B 2.86 5.71 12.86 2.14 2.99 14.30 12.12 13.64 18.94 11.11 11.11 19.44 3.17 4.76 87.30 0.00 0.00 17.46
qwen2.5-coder-instruct 32B 12.86 15.71 30.00 4.07 4.96 18.40 9.85 15.15 13.64 25.00 27.78 30.56 26.98 33.33 96.83 4.76 4.76 17.46
deepseek-coder-instruct 6.7B 4.29 8.57 4.29 1.73 2.82 8.79 3.79 5.30 17.42 5.56 5.56 19.44 0.00 4.76 60.32 0.00 0.00 9.52
deepseek-coder-instruct 33B 2.86 2.86 18.57 1.94 3.23 11.27 4.55 6.82 22.73 11.11 13.89 22.22 3.17 25.40 92.06 0.00 1.59 15.87
Language-wise Average 12.20 6.15 11.99 15.97 34.92 6.35
Average 7.86 10.54 18.21 2.53 3.39 12.53 8.71 11.27 16.00 12.85 14.24 20.83 10.91 17.86 75.99 1.79 2.38 14.88
[…]
* The improvement ratios shown in “Pass@5” column are calculated by comparing with the scores in Pass@1, and the ratios shown in “Few-shot” column are calculated by comparing with the scores in Pass@5.
* ACSL and ACSL-WP: ‘ACSL’ uses ‘frama-c’ tool for syntax and basic ACSL checking; ‘ACSL-WP’ uses ‘frama-c’ with the WP plugin for specification verification.
Table 5: RQ2: Language-wise LLMs’ Performance. Pass@1 and Pass@5 Accuracy in Generating Proof Segments Across Formal Specification Languages Under Zero/Few-shot without fine-tuning.
MATH Reasoning Coding
Fine-tuning Dataset MATH GSM-8K Average bbh HumanEval MBPP Average Average
UltraChat 17.54 61.33 39.44 62.64 19.51 36.4 27.96 39.48
[…]
tulu3 27.36 75.82 51.59 62.47 64.02 48.8 56.41 55.01
tulu3 + fma 29.48 7.75% $\uparrow$ 75.44 $0\downarrow 5$2.46 1.69% $\uparrow$ 63.16 1.10% $\uparrow$ 64.63 0.95% $\uparrow$ 49.4 1.23% $\uparrow$ 57.02 1.07% $\uparrow$ 55.76 1.37% $\uparrow$
Table 6: RQ4: Capability Migration from fm-alpaca to Math, Reasoning, and Coding.
4.3 RQ2. Formal Specification Languages-wise Capability
Table 5 shows the LLMs’ performance across formal specification languages in the task of generating proof segments (SegGen). This task accounts for the most instructions and serves as the basic capability for other proof generation tasks. We can see that LLMs perform the best in ACSL (average: 34.92%), followed by Dafny (15.97%) while performing unsatisfactorily in other formal specification languages. The observation is reasonable because the syntax of ACSL is basically an annotation of C language, while Dafny shares similar grammar as C# and Java. Thus, generating proof segments in ACSL and Dafny is generally easier than generating other specification languages.
Note that though LLMs are generally good at generating syntactically correct and semantically consistent (i.e., program behaviors are consistent with the annotated ACSL specifications) according to the ‘ACSL’ column in Table 3, the specifications are usually insufficient for the verification according to the ‘ACSL-WP’ column in Table 3. The results in the two ACSL columns indicate that there is a large room to improve for generating not only syntactically/semantically correct specifications, but also that the generated specifications need to be sufficient to verify the properties under verification.
In addition, we explore whether increasing the attempts (1 $\rightarrow$ 5) with a higher temperature (0.0 $\rightarrow$ 0.2) and in-context learning could bring about improvement. The improvement ratios are shown in red in Table 5. The results of Pass@5 are better than those of Pass@1, with an average score increase from 10.82% (Dafny) to 63.64% (ACSL) in different languages. Moreover, when using 3-shot, the performance increases dramatically, with 51.33% (Dafny) to over five times (ACSL) improvement compared with zero-shot Pass@5. The results indicate the potential of in-context learning in generating correct specification languages.
4.4 RQ3. Improvement by Fine-tuning
[…]
Yet, drops can be observed in fill-in-the-middle tasks (i.e., ProofInfill and Code2Proof). The results echo the observation made in RQ1 (Section 4.2), where the large LLMs perform worse than small LLMs in fill-in-the-middle tasks. The results also indicate the necessity of adopting different fine-tuning strategies other than instruction tuning only.
Fine-tuning Datasets: Take a closer look at the LLMs fine-tuned with general-purpose datasets (i.e., llama3.1-ultrachat and llama3.1-tulu) in Table 3, with them only, no or opposite effects can be observed. The results indicate the complementarity of fm-alpaca and existing general-purpose fine-tuning datasets. Additionally, by combining with general-purpose datasets, the performance can be further improved (e.g., llama3.1-tulu-fma).
Comparison with Few-shot: Compared with the best results in Table 5 achieved by 3-shot, the results after fine-tuning (Table 3) still generally outperform the 3-shot results. The results indicate that although in-context learning can improve LLMs’ performance, the enhancement is limited. Further significant improvements still require fine-tuning with formal data. This may also suggest that in-context learning alone cannot adequately address capability deficits in formal verification tasks but rather stem from a lack of knowledge.
4.5 RQ4. Capability Migration from Formal Verification to Related Tasks
Finally, we explore whether fine-tuning with fm-alpaca could benefit related capabilities. Table 6 shows the results. The base model is llama3.1-8B, fine-tuned under two base fine-tuned datasets with and without fm-alpaca. On average, with fm-alpaca, an increase of 1.37% to 5.15% can be observed. Interestingly, a dramatic increase (62.53%) can be observed in HumanEval compared with the performance of the model that is only fine-tuned with UltraChat. The experiment may indicate that feeding more formal data may improve LLMs’ coding, reasoning, and math capabilities.
5 Conclusion
This paper contributes a comprehensive assessment and formulation to understand LLMs’ capability in formal verification. We constructed 18k high-quality instruction-response pairs across five formal specification languages in six tasks. The fine-tuned models, fine-tuning data, and the benchmark are released to facilitate subsequent studies.
Limitations
This paper has two primary limitations that offer avenues for future research. First, the primary limitation of our work is that our benchmark relies on model-generated data. While this approach effectively reduces manual efforts; it may introduce biases and data leakage issues in the dataset towards the models that generated the data. To address this limitation, we use gpt-4o to generate the natural language descriptions, while during the evaluation, we use other LLMs for evaluation. Second, another limitation of our work lies in the validation design. When creating ProofInfill and ProofComplete data, it is possible that the properties to be verified or theorems to be proven are masked. If LLMs happened not to generate these properties/theorems, the generated “proofs/models” could escape the verifier/checker, mistakenly labeling the output as correct. To avoid this scenario, we include the requirement descriptions as part of the input, guiding LLMs to generate the necessary properties or theorems without omission.
6 Acknowledgments
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported in part by the Hong Kong Research Grants Council/General Research Fund (HKSAR RGC/GRF, Grant No. 16206524), the National Natural Science Foundation of China (Grant Nos. 62372193, 62192734, 62302375, 62472339), CAS Project for Young Scientists in Basic Research (Grant No.YSBR-040) and the Basic Research Program of ISCAS (Grant No. ISCAS-ZD-202402).
Appendix A Related Work
The formal specification datasets or benchmarks offer a standard, well-defined set of problems, providing a shared challenge that helps build a community of practice among researchers. According to different verification techniques, the existing benchmarks mainly fall into two categories; we discuss them separately.
A.1 Theorem Proving Datasets
Formal theorem proving represents theorems and proofs in a machine-verifiable format Cook (2023), ensuring their correctness using rigorous logical rules. A recent survey Li et al. (2024) summarized the existing datasets for theorem proving. In particular, the informal benchmarks craft the proofs from various sources such as ProofWiki, textbooks, and public corpus. NL-PS Ferreira and Freitas (2020) first builds a natural language premise selection dataset source from ProofWiki. Similarly, NaturalProofs Welleck et al. (2021) further incorporates data from Stacks and textbooks, resulting in a dataset with roughly 25k examples. Adapted from it, NaturalProofs-Gen Welleck et al. (2022) contains around 14.5k theorems for informal proof generation. Moreover, MATcH Li et al. (2023) constructs over 180k statement-proof pairs for matching using the MREC corpus  [^2] https://mir.fi.muni.cz/MREC/.
For formal datasets, a line of efforts focuses on extracting and cleaning theorems and proofs written in various specification languages (e.g., Coq, Isabelle, Lean) from established formal libraries and verification projects. For example, LeanDojo Yang et al. (2023) extracts over 98k theorems and proofs with 130k premises from Lean mathlib mathlib Community (2020). Besides extracting data from existing projects, several works manually annotate or formalize the problems in natural language. For example, MiniF2F Zheng et al. (2021) manually formalizes 488 Olympiad-level problems across 4 proof systems and equally splits them into a validation set and a test set. FIMO Liu et al. (2023) and ProofNet Azerbayev et al. (2023) formalize the theorem statements of the International Mathematical Olympiad and undergraduate-level problems in Lean. In addition, datasets for Dafny also attract research contributions because industries like Amazon adopted Dafny to verify cryptographic libraries, authorization protocols, a random number generator, and the Ethereum virtual machine. Dafny datasets such as CloverBench Sun et al. (2024) and DafnyGym Mugnier et al. (2024).
A.2 Model checking datasets
Model checking is an automated technique used in computer science and formal methods to verify the correctness of systems, particularly those with finite state spaces. It systematically checks whether a system’s model satisfies a given specification, usually expressed in formal specification languages. The basic idea is to explore all possible system states to ensure the desired properties hold in every conceivable scenario.
Model checking benchmarks are less than that for theorem proving. Currently, there are few model-checking benchmarks for proving, while several model-checking subjects are going with specific model-checking languages such as CMurphi Della Penna et al. (2013) and TLA+ Yu et al. (1999). In particular, CMurphi is a software tool used to verify concurrent and distributed systems through explicit state enumeration. It implements the Murphi verification language, which allows users to describe finite-state systems in a procedural style. The core principle behind CMurphi is to explore the state space of a system exhaustively to check for violations of specified invariants or properties. Another example is TLA+ (Temporal Logic of Actions), a high-level language for modeling programs and systems suitable for concurrent and distributed systems.
Appendix B Proportion of Failures Caused by Syntax Error
We listed the proportions of failures caused by syntax errors for each LLM and each task in Table 8. We used a set of pre-defined keywords (summarized in Table 7 to identify if a verification failure is caused by syntax errors. Specifically, we consider a failure caused by syntax error if its error message contains at least one keyword in Table 7.
Language Keywords
1 Coq “Syntax Error:”
2 Lean4 “unexpected token ”, “unknown identifier”, “type mismatch”
3 ACSL “unexpected token”
4 Dafny “type errors detected”
5 TLA+ “***Parse Error***”, “Unknown operator”
Table 7: Keywords for identifying syntax error raised by each language’s verifier.
LLMs Size ProofGen SegGen ProofComplete ProofInfill Cd2Prf
TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL
llama3.1-instruct 8B 0/13 18/45 34/58 0/10 1/14 43/69 742/2910 78/124 4/33 4/63 4/11 45/81 16/34 1/5 16/34 44/157 27/60 0/13 1/14 0/14
llama3.1-instruct 70B 4/12 10/44 40/53 1/8 1/14 43/53 71/2802 75/114 16/32 5/61 6/8 44/81 10/31 1/5 15/34 39/153 22/49 0/13 2/14 3/14
qwen2.5-instruct 7B 12/13 13/46 33/56 1/10 0/14 26/69 302/2914 53/122 5/34 6/62 9/11 35/73 8/29 1/5 14/34 66/156 31/53 0/13 4/14 3/14
qwen2.5-instruct 72B 10/12 6/44 28/48 1/8 9/13 47/61 40/2848 54/116 11/27 0/60 7/11 35/74 8/24 1/5 13/34 59/149 20/49 0/11 0/14 1/14
qwen2.5-coder-instruct 7B 11/13 18/46 38/55 0/10 0/14 60/68 22/2882 72/116 6/32 53/63 10/11 12/80 8/32 2/6 17/34 34/159 26/50 1/12 0/14 0/14
qwen2.5-coder-instruct 32B 11/13 12/44 29/52 1/10 2/14 50/61 31/2825 68/119 9/27 25/60 5/8 24/79 6/23 3/4 17/33 51/152 16/43 0/12 0/12 0/12
deepseek-coder-instruct 7B 10/13 4/46 31/54 1/10 0/14 35/67 9/2894 62/127 3/34 43/63 9/10 14/82 11/31 2/5 22/34 18/160 23/56 0/10 1/14 1/14
deepseek-coder-instruct 33B 12/13 0/47 46/58 4/10 0/14 66/68 52/2888 81/126 8/32 45/63 10/11 3/81 6/31 2/4 19/34 6/161 23/50 0/14 0/14 1/14
star-coder-instruct 15B 8/12 0/46 27/54 2/7 3/14 36/51 10/2887 67/124 12/26 1/63 6/8 4/74 7/27 1/4 20/34 6/149 23/49 0/12 0/14 1/14
  • 1
    Denominator represents the total number of failures.
Table 8: Proportion of Verification Failures Caused by Syntax Errors
Appendix C Example Specifications in fm-alpaca and fm-bench
The examples of the five formal specification languages are shown in Figure 4.
Figure 4: Formal Specification Languages in FM-bench
Appendix D Collected Repositories
We listed the repositories that were collected for data construction in the following. Note that one can easily add more repositories into fm-alpaca and fm-bench.
Appendix E Complete Evaluation Result
The Pass@1 and Pass@5 are shown in Table 9. It is a completed version of Table 3.
Pass @ 1
ProofGen SegGen ProofComplete ProofInfill Cd2Prf
LLMs Size TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL-WP ACSL ACSL-WP
W/o Fine-tuning
llama3.1-instruct 8B 0.00 4.26 1.69 0.00 14.29 0.00 1.43 1.19 6.06 8.33 0.00 0.00 0.00 3.57 5.56 28.57 0.00 4.85 7.69 13.33 21.43 0.00 57.14 0.00
llama3.1-instruct 70B 7.69 6.38 10.17 20.00 7.14 0.00 24.29 4.86 13.64 11.11 39.68 3.17 27.27 3.57 13.89 28.57 0.00 7.27 24.62 13.33 21.43 0.00 21.43 0.00
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 21.43 0.00 1.43 1.05 7.58 5.56 1.59 1.59 0.00 13.10 19.44 28.57 0.00 5.45 18.46 13.33 21.43 0.00 14.29 0.00
qwen2.5-instruct 72B 7.69 6.38 18.64 20.00 21.43 7.14 12.86 3.29 12.12 25.00 12.70 4.76 0.00 11.90 33.33 28.57 0.00 9.70 24.62 26.67 35.71 0.00 57.14 0.00
qwen2.5-coder-instruct 7B 0.00 2.13 6.78 0.00 0.00 0.00 2.86 2.14 12.12 11.11 3.17 0.00 0.00 4.76 11.11 14.29 0.00 3.64 23.08 20.00 0.00 0.00 0.00 0.00
qwen2.5-coder-instruct 32B 0.00 6.38 11.86 0.00 21.43 0.00 12.86 4.07 9.85 25.00 26.98 4.76 27.27 5.95 36.11 42.86 2.94 7.88 33.85 20.00 71.43 14.29 57.14 14.29
star-coder-instruct 15B 7.69 2.13 8.47 30.00 28.57 0.00 27.14 1.97 6.06 27.78 0.00 0.00 27.27 11.90 25.00 42.86 0.00 9.70 24.62 20.00 35.71 0.00 28.57 0.00
deepseek-coder-instruct 7B 0.00 2.13 8.47 0.00 0.00 0.00 4.29 1.73 3.79 5.56 0.00 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 0.00 0.00 0.00 0.00
deepseek-coder-instruct 33B 0.00 0.00 1.69 0.00 0.00 0.00 2.86 1.94 4.55 11.11 3.17 0.00 0.00 3.57 13.89 42.86 0.00 2.42 23.08 6.67 7.14 0.00 7.14 0.00
deepseek-r1 xxB 30.77 8.51 30.51 30.00 35.71 7.14 22.86 10.70 21.21 22.22 49.21 3.17 45.45 15.48 33.33 28.57 5.88 9.70 24.62 6.67 42.86 21.43 64.29 14.29
W/ Fine-tuning
llama3.1-sft 8B 0.00 6.38 8.47 20.00 0.00 0.00 38.57 29.58 21.97 25.00 88.89 17.46 36.36 8.33 11.11 28.57 0.00 0.00 7.69 0.00 21.43 0.00 21.43 0.00
llama3.1-ultrachat 8B 0.00 6.38 3.39 0.00 0.00 0.00 0.00 0.00 3.79 0.00 0.00 0.00 0.00 5.95 13.89 14.29 0.00 4.85 9.23 0.00 14.29 7.14 0.00 0.00
llama3.1-ultrachat-fm 8B 0.00 8.51 10.17 30.00 0.00 0.00 41.43 35.79 29.55 33.33 95.24 17.46 18.18 8.33 11.11 28.57 2.94 0.00 7.69 0.00 28.57 0.00 35.71 0.00
llama3.1-tulu 8B 0.00 2.13 1.69 0.00 0.00 0.00 0.00 0.71 6.06 0.00 0.00 0.00 0.00 5.95 8.33 28.57 0.00 1.82 10.77 0.00 7.14 0.00 14.29 0.00
llama3.1-tulu-fm 8B 0.00 4.26 11.86 30.00 7.14 0.00 42.86 36.64 27.27 36.11 98.41 17.46 18.18 10.71 11.11 42.86 2.94 2.42 6.15 0.00 35.71 0.00 50.00 0.00
qwen2.5-sft 7B 0.00 4.26 11.86 10.00 0.00 0.00 38.57 27.95 27.27 22.22 87.30 17.46 27.27 9.52 13.89 28.57 0.00 0.00 7.69 0.00 21.43 0.00 14.29 0.00
qwen2.5-coder-sft 7B 0.00 6.38 18.64 20.00 7.14 0.00 44.29 36.50 34.09 33.33 98.41 17.46 36.36 9.52 16.67 28.57 5.88 1.82 13.85 13.33 42.86 0.00 42.86 0.00
deepseek-coder-sft 7B 0.00 2.13 16.95 0.00 0.00 0.00 34.29 25.30 31.06 25.00 84.13 15.87 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 14.29 7.14 14.29 7.14
Pass @ 5
ProofGen SegGen ProofComplete ProofInfill Cd2Prf
LLMs Size TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP ACSL ACSL-WP
W/o Fine-tuning
llama3.1-instruct 8B 0.00 2.13 5.08 0.00 35.71 0.00 4.29 1.90 8.33 11.11 3.17 0.00 0.00 3.57 8.33 28.57 0.00 7.88 20.00 20.00 42.86 0.00 78.57 7.14
llama3.1-instruct 70B 7.69 8.51 13.56 20.00 35.71 21.43 27.14 6.01 17.42 13.89 52.38 4.76 36.36 11.90 27.78 28.57 0.00 10.30 30.77 20.00 35.71 7.14 50.00 14.29
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 28.57 0.00 2.86 1.39 9.09 5.56 3.17 3.17 0.00 16.67 22.22 28.57 0.00 5.45 20.00 20.00 42.86 0.00 42.86 0.00
qwen2.5-instruct 72B 15.38 6.38 22.03 30.00 21.43 7.14 17.14 3.80 14.39 25.00 15.87 4.76 9.09 16.67 38.89 28.57 0.00 11.52 26.15 33.33 57.14 14.29 92.86 7.14
qwen2.5-coder-instruct 7B 0.00 4.26 10.17 0.00 0.00 0.00 5.71 2.99 13.64 11.11 4.76 0.00 0.00 11.90 19.44 28.57 0.00 4.24 30.77 20.00 7.14 0.00 0.00 0.00
qwen2.5-coder-instruct 32B 7.69 6.38 15.25 10.00 28.57 0.00 15.71 4.96 15.15 27.78 33.33 4.76 36.36 10.71 47.22 42.86 2.94 10.91 33.85 20.00 92.86 28.57 71.43 28.57
star-coder-instruct 15B 15.38 4.26 10.17 30.00 42.86 0.00 34.29 3.46 7.58 27.78 0.00 0.00 45.45 16.67 41.67 57.14 0.00 16.97 38.46 26.67 42.86 0.00 42.86 7.14
deepseek-coder-instruct 7B 7.69 2.13 10.17 0.00 0.00 0.00 8.57 2.82 5.30 5.56 4.76 0.00 18.18 9.52 16.67 28.57 0.00 4.24 18.46 33.33 21.43 0.00 21.43 0.00
deepseek-coder-instruct 33B 0.00 6.38 6.78 20.00 7.14 0.00 2.86 3.23 6.82 13.89 25.40 1.59 0.00 11.90 19.44 42.86 0.00 10.30 40.00 26.67 14.29 7.14 14.29 0.00
W/ Fine-tuning
llama3.1-sft 8B 0.00 6.38 15.25 20.00 0.00 0.00 41.43 34.57 26.52 27.78 90.48 17.46 36.36 9.52 11.11 28.57 2.94 0.00 13.85 0.00 35.71 0.00 57.14 0.00
llama3.1-ultrachat 8B 0.00 8.51 5.08 0.00 0.00 0.00 0.00 0.07 7.58 0.00 1.59 0.00 0.00 4.76 22.22 14.29 0.00 4.85 9.23 6.67 14.29 7.14 0.00 0.00
llama3.1-ultrachat-fm 8B 0.00 10.64 15.25 30.00 7.14 0.00 42.86 39.86 31.06 38.89 96.83 17.46 18.18 9.52 13.89 28.57 11.76 0.00 10.77 0.00 42.86 7.14 64.29 7.14
llama3.1-tulu 8B 0.00 6.38 1.69 0.00 0.00 0.00 0.00 1.43 6.06 0.00 3.17 0.00 0.00 13.10 30.56 28.57 0.00 6.67 21.54 13.33 21.43 7.14 21.43 0.00
llama3.1-tulu-fm 8B 0.00 6.38 11.86 30.00 7.14 0.00 44.29 40.20 32.58 38.89 98.41 17.46 18.18 13.10 11.11 42.86 2.94 4.24 9.23 0.00 71.43 7.14 64.29 0.00
qwen2.5-sft 7B 0.00 4.26 13.56 10.00 7.14 0.00 40.00 32.77 30.30 27.78 88.89 17.46 45.45 10.71 19.44 28.57 2.94 1.21 12.31 0.00 42.86 0.00 35.71 0.00
qwen2.5-coder-sft 7B 0.00 8.51 22.03 30.00 7.14 0.00 44.29 40.81 35.61 36.11 98.41 17.46 54.55 10.71 19.44 42.86 8.82 4.24 20.00 13.33 64.29 0.00 57.14 0.00
deepseek-coder-sft 7B 0.00 4.26 16.95 0.00 0.00 0.00 35.71 30.70 36.36 25.00 85.71 15.87 54.55 13.10 16.67 28.57 2.94 4.24 9.23 6.67 14.29 7.14 28.57 7.14
Table 9: Full Experiment Results on Pass@1 and Pass@5
Appendix F Prompt Design
We listed the prompts that are used for data preparation and inference in the following. For data preparation, as shown in Figure 5, to generate descriptions for the given proof segments, the prompt template consists of five parts: (1) Role description, (2) Domain knowledge of TLA+, (3) Task description, (4) Few-shot examples (we show one example in the figure, while three-shots were used in RQ2), (5) The proof or proof segment to be summarized.
[…]
## Code
Return(c,S) ==
[…]
<A proof segment>
Figure 5: Prompt for generating TLA+ description. The prompt templates for other formal specification languages are in the same structure.
Prompt for SegGen Task (lang: a placeholder to be replaced by each formal specification language name.) # Task Description (SegGen) 1. Translate the given natural language into {lang} syntax. 2. Model the intention written in natural language using {lang}. 3. Express the requirement using {lang}. 4. Model the given natural language into {lang}. 5. Translate the given requirement using {lang}’s syntax and semantics. (Randomly choose one of the above.) You only need to return the {lang} formal specification without explanation. # Input
<Input goes here>
Figure 6: Prompt for SegGen Task.
Prompt for ProofGen Task (lang: a placeholder to be replaced by each formal specification language name.) # Task Description (ProofGen) 1. Translate the given requirements into lang syntax. 2. Model the given requirements written in natural language using lang. 3. Express the requirements using lang. 4. Model the given requirements written in natural language into lang. 5. Translate the given requirements into lang’s syntax and semantics. (Randomly choose one of the above.) (For ACSL): You only need to return the lang formal specification with the code without explanation. (For others): You only need to return the lang formal specification without explanation. # Input
<Input goes here>
Figure 7: Prompt for ProofGen Task.
Prompt for ProofComplete Task (lang: a placeholder to be replaced by each formal specification language name.) # Task Description (ProofComplete) 1. Please complete the following formal proof in formal specification language lang according to the given requirement. 2. Please complete the following formal proof in lang according to the given requirement. 3. Please complete the given formal proof in lang following the requirement below. 4. Please complete the following formal proof in lang according to the requirement below. 5. Please complete the following formal proof in lang according to the given requirement. (Randomly choose one of the above.) You only need to return the completed lang formal specification (together with the provided formal specification) without explanation. # Input
<Input goes here>
Figure 8: Prompt for ProofComplete Task.
Prompt for ProofInfill Task (lang: a placeholder to be replaced by each formal specification language name.) # Task Description (ProofInfill) 1. Please fill in the [MASK] in the following formal proof in formal specification language lang according to the given requirement. 2. Please fill in the [MASK] in the following formal proof in lang according to the given requirement. 3. Please complete the given formal proof in lang following the requirement below by filling in the [MASK]. 4. Please fill in the [MASK] in the following formal proof in lang according to the requirements below. 5. Please fill in the [MASK] in the following formal proof in lang according to the given requirement. (Randomly choose one of the above.) You only need to return the completed lang formal specification (together with the provided formal specification) without explanation. # Input
<Input goes here>
Figure 9: Prompt for ProofInfill Task.
Prompt for Code2Proof Task # Task Description (Code2Proof) 1. Please fill in the [MASK] in ACSL according to the given requirement and ACSL specification. 2. Please fill in the [MASK] in ACSL according to the given requirement. 3. Please fill in the [MASK] in ACSL according to the given ACSL specification. 4. Please fill in the [MASK] in ACSL according to the given requirement and ACSL specification. 5. Please infill the [MASK] in ACSL according to the given requirement. (Randomly choose one of the above.) You only need to return the completed ACSL formal specification (together with the provided formal specifications and C programs) without explanation. # Input
<Input goes here>
Figure 10: Prompt for Code2Proof Task.